OpenStudy (anonymous):

Tell me If I am right Find the x and y intercepts f(x)=-x^2+7x+30 x =(3,10) y=(0,30)

OpenStudy (anonymous):

yeah

OpenStudy (saifoo.khan):

Wrong^

OpenStudy (anonymous):

ok the y intercept is (0, 3) because if you replace x by 0 you get y = 30

OpenStudy (saifoo.khan):

ok the y intercept is (0, 30) because if you replace x by 0 you get y = 30

OpenStudy (anonymous):

Your y-int is good. To find the x-ints (there are two) let y=o and solve for x -x^2+7x+30=0 x^2-7x-30=0 multi through by -1 (x-10)(x+3)=0 factor x-10=0 or x+3=0 zero-factor theorem x=10 or x=-3 points of intersection on x-axis (-3,0) and (10,0)

OpenStudy (saifoo.khan):

correct^

OpenStudy (anonymous):

the x intercept should look like \[(a,0)\] and you should have two of them

OpenStudy (anonymous):

server is beginning to drive me nuts

OpenStudy (anonymous):

i agree!

OpenStudy (anonymous):

Shoot, I'm sorry, there I deleted my wrong answer, ^.^

OpenStudy (anonymous):

one is right. but you should write \[(10,0)\] as one of them. the other is not 3, but rather -3, so your answer should look like \[(10,0), (-3,0)\]

OpenStudy (anonymous):

so x is (10,-3)?

OpenStudy (saifoo.khan):

NO.

OpenStudy (saifoo.khan):

(-3,0) and (10,0)

OpenStudy (anonymous):

first coordinate is the x value, second coordinate is 0

OpenStudy (anonymous):

there are two intercepts for x

OpenStudy (anonymous):

all points on the x-axis have y-coordinate 0